A truck driver is attempting to deliver some furniture.first he travels 8km east and then he turns around and travels 3 km West finally he turns again and travels 12 km east to his destination
1
Expert's answer
2019-11-21T08:51:34-0500

Total distance is


"d=8+3+12=23\ km"

Total displacement is


"s=8-3+12=17\ km\ to\ the\ east"